OpenStudy (anonymous):

Explain the steps used to solve x - 2 = v(x + 10) and check for extraneous solutions.

OpenStudy (anonymous):

what's the v for?

OpenStudy (anonymous):

i dont know.....thats what it says

OpenStudy (anonymous):

well x - 2 = v(x+10) x - 2 = x*v + 10v x - x*v = 2 + 10v x(1-v) = 2+10v x =(2 + 10v)/(1-v) in this case, your answer would depend on the value of v. notice however, that v can't = 1. if it does, then you are dividing by zero. so the solutions are all values of v except v = 1